Thursday 01/10/13

8:30A – 10:30P

  • Flip collection plate 8:30A
  • Fixing embryos 10:30A.  (20 min 8% FA reaction)
  • O/N confocal runs — 700 at speed 4 still running.  5 not quite as clean.
  • copy data.
  • Make new fly plates (~4 tubes of plates).
  • working on Fig 1 layout for new version of bistability ms.

Cell Culture

  • finish DFISH staining for DFISH second.  Imaging these wells on STORM2


  • Align back-optics
  • bead imaging (just 3D calibration.  need to take chromatic beads tomorrow).  Should shoot another set of 3D beads with new, less dense bead slide.
  • Need to check laser powers!  started imaging with too low laser power the K9me2.  Should repeat tomorrow

Probe making

  • make fresh 10% APS solution
  • denaturing gel cast 1: polymerizes immediately upon addition of APS.  
  • Gel 2 cooled down to RT.  Too much H2O evaporated, (should cover with parafilm to minimize evaporation).  Urea crashed out of solution.
  • Re-supended urea with gently heating.  Added 1/2 x APS.
  • however not enough gel to fill up to comb left.  Gel fully polymerized within 1 min of pouring into cast.
  • Resuspending concentrated probe.  Will run out on new gel tomorrow.


This entry was posted in Summaries and tagged , , . Bookmark the permalink.